OpenStudy (anonymous):

can someone help me with this...plzzzzzzz! solve the equation below for x. 9(x - 5) = 4x - 5

OpenStudy (anonymous):

I think I can help you @kimberly_r Please give me a few minutes to solve it.

OpenStudy (anonymous):

yaaay ok thx

OpenStudy (anonymous):

So what you get automatically is that 9 = 4x, do you understand so far?

OpenStudy (anonymous):

9x-15=4x-5 right??

OpenStudy (anonymous):

i hateeeeeeeeee math :(

OpenStudy (anonymous):

Yes! Good job. 9x - 45 = 4x - 5 So, we can cancel -5 by adding 5 to it. Whatever we do to that side must be done to the other side. So now we have 9x - 50 = 4x

OpenStudy (anonymous):

Do you understand that part.

OpenStudy (anonymous):

yeaaaa but how do i get rid of the 9x now?

OpenStudy (anonymous):

Personally I got stuck here too. What I did was realize that 9-5=4 So 50/5 = 10 So it seems like x=10 Please correct me if I am wrong.

OpenStudy (radar):

You may be wrong, because 9(10-5) = 40 - 5 45 = 35 is not correctt.

OpenStudy (radar):

Start from where you had 9x-45 = 4x -5 Try adding 5 to both sides again. That is where you erred.

OpenStudy (anonymous):

awww i think i figured it out

OpenStudy (anonymous):

is it x=8?

OpenStudy (radar):

Yes

OpenStudy (radar):

Good luck with your studies.

OpenStudy (anonymous):

yaaay thx